OpenStudy (anonymous):

Why would one use the square root property of equality instead of factoring to solve quadratic equations?

OpenStudy (anonymous):

some quadratic equations cannot be factored

OpenStudy (anonymous):

i meant factored easily. What I mean is factoring works in a very tiny percentage of quadratic equations. in most equations it is not evident what the factors of a particular equation are.

OpenStudy (anonymous):

example would be \[x^2+6x-1=0\] which does not factor over integers. so you would say \[x^2+6x=1\] \[(x+3)^2=1+9=10\] \[x+3=\pm\sqrt{10}\] \[x=-3\pm\sqrt{10}\] because it does not factor using integers you use square root property

OpenStudy (anonymous):

Thank you both so much. You were a big help. Now it makes sense